When an object is thrown vertically upward it raises to maximum height of 30m. How high is the object when it has lost one third of its kinetic energy? What is the speed of the object at this point?


1
Expert's answer
2021-05-06T07:41:45-0400

"mgH=mv^2\/2"


"\frac{1}{3}KE=\frac{1}{3}\cdot\frac{mv^2}{2}=\frac{1}{3}\cdot mgH"


"\frac{1}{3}\cdot mgH=mgh\to h=H/3=30/3=10\ (m)" . Answer


"mgH-\frac{1}{3}\cdot mgH=\frac{2}{3}\cdot mgH"


"\frac{2}{3}\cdot mgH=\frac{mv^2}{2}\to v=\sqrt{\frac{4gH}{3}}=\sqrt{\frac{4\cdot9.81\cdot 30}{3}}=19.8\ (m/s)" . Answer